# NAMET: Robust Massive Model Editing via Noise-Aware Memory Optimization

### v1

@@ 4.2 Empirical Study of Embedding Collision
Model Prefix Lens Editing Acc. low r-k Acc. top r-k Acc. low k-k Acc. top k-k Acc. low k-k-$\lambda$ Acc. top k-k-$\lambda$ Acc.
LLaMA2-7B 0 98.50% - - - - - -
5 84.15% 83.00% 47.00% 73.00% 39.00% 93.00% 27.00%
10 80.35% 79.00% 44.00% 68.00% 34.00% 90.00% 24.00%
50 77.40% 70.00% 37.00% 54.00% 28.00% 90.00% 21.00%
Deepseek-7B 0 99.50% - - - - - -
5 79.35% 82.00% 36.00% 88.00% 31.00% 71.00% 58.00%
10 84.25% 92.00% 54.00% 97.00% 38.00% 85.00% 58.00%
50 85.10% 91.00% 47.00% 90.00% 37.00% 91.00% 62.00%
@@
Empirical Verification of Hypothesis 2. We proceed by editing 200 facts from the CounterFact dataset in LLaMA2-7B and Deepseek-7B using MEMIT. To evaluate each edit, we measure the average editing accuracy when the original factual association is preceded by prefixes of varying lengths. Specifically, we vary the prefix length from 5 to 50 tokens and report the average editing accuracy across all edited items. Additionally, we compute the average editing accuracy of the 10 items with the highest collision scores for r-k, k-k, and k-k-$\lambda$ pairs separately. We also evaluate the 10 items with the lowest collision scores for the same types of pairs, in order to illustrate how embedding collisions influence the robustness of edited knowledge under long-prefix conditions.
@@
Table 2 demonstrates the degradation of editing performance when evaluating edited facts with prefixes. For LLaMA2-7B, the accuracy drops from 98.5% to 84.15% when preceded by 5 tokens, and further to 77.40% for 50 tokens. The degradation phenomenon is also observed for Deepseek-7B, where the editing accuracy drops around 15% when preceded by 50-token prefixes. The robustness to prefix perturbations varies significantly between items that are prone to embedding collisions and those that are not. Corresponding to Hypothesis 2, items with lower collision degrees exhibit editing accuracy comparable to the overall average, whereas items with higher collision degrees suffer a substantial decline, with accuracy dropping to approximately 30%. This discrepancy is especially pronounced when collision is measured using r-k pairs. For both models, items with low collision degrees achieve editing accuracy around 80%, exceeding that of collision-prone items by over 40%.
These results highlight connections between embedding collisions, and the effectiveness and robustness of massive editing. Since key embeddings are fixed by the model and residuals are derived from target memory embeddings, advanced ME methods can only improve the isolatedness of memory embeddings across different knowledge items.
5 Adjustment to Existing Model Editing Method and Metric
Figure 4: (LEFT) Illustration of changing padding strategy. (RIGHT) Overview and information flow of NAMET.
Motivated by the above results, we propose NAMET, an effective one-line modification to MEMIT, which generates more isolated memory embeddings through introducing noise during optimization. This improves both the model’s capacity to edit multiple facts concurrently under practical metric, and the robustness when evaluating the edited items with prefixes. We then elaborate on the details.
Mitigating Embedding Collisions via Noise Injection. Padding tokens, which carry no semantic content, are ideal for injecting noise during optimizing target memory embeddings. Based on this insight, we propose to change the padding strategy from right-padding to left-padding, thereby prepending [unk] tokens before factual prompts during memory optimization. Formally, let $h_{i}^{L}$ denote the hidden state at layer $L$ corresponding to the $i$-th target factual association (subject $s_{i}$, relation $r_{i}$, object $o_{i}$). We optimize upon it to obtain target hidden state $z_{i}$ via:
$z_{i}=h_{i}^{L}+\underset{\delta_{i}}{\arg\min}\frac{1}{N_{FP}}\sum_{j=1}^{N_{FP}}-\log\mathds{P}_{G(h_{i}^{L}+=\delta_{i})}[o_{i}\mid f_{j}\oplus tp(s_{i},r_{i})]$ (10)
Here, $\delta_{i}$ is optimized to maximize the model’s confidence in predicting the target object $o_{i}$. $f_{j}\oplus tp(s_{i},r_{i})$ concatenates random prefixes $f_{j}$ with templated prompt $tp(s_{i},r_{i})$ for each target edit $i$. $G_{(h_{i}^{L}+=\delta_{i})}$ denotes that we add change $\delta_{i}$ to the hidden state $h_{i}^{L}$ during the model execution.
As illustrated in Figure 4, each edit is augmented with $N_{FP}-1$ prefixes of fixed length, resulting in $N_{FP}$ factual prompts per edit. The first prompt contains no prefix and is thus shorter, causing the tokenizer to prepend [unk] tokens under left-padding. These padding tokens introduce noise in hidden states, promoting more isolated memory representations during optimization and mitigating embedding collisions across knowledge items. Notably, this method requires only a one-line modification to the original MEMIT implementation. The complete algorithm is provided in Appendix C.
Practical Evaluation Metric. We propose to evaluate the editing performance of ME techniques by requiring the edited models to strictly examine whether explicit target objects are retrieved, as demonstrated in Figure 1. The editing efficacy is then defined as:
@@ 6 Experiments
In this section, we empirically focus on evaluating the following research questions (RQs), while additional experiments are provided in Appendix D. Codes are available at https://github.com/ybdai7/NAMET-massive-editing.
  • RQ1. Can NAMET generate more isolated embeddings for different knowledge items?
  • RQ2. How does NAMET perform on massive editing tasks compared with baselines for various LLMs? Can it excel under the practical metric?
  • RQ3. How does NAMET perform when evaluating edited facts with prefixes?
@@ 6.1 Experiments Setup
Models, datasets, and baselines. We conduct extensive experiments on various LLMs, including LLaMA2-7B [30], LLaMA2-13B [30], Falcon-7B [2], Qwen-2.5-7B [34], Deepseek-base-7B [3], and LLaMA3-8B [31]. We consider a range of ME techniques as baselines: FT [39], MEND [21], ROME [19], MEMIT [20], PMET [15], and ALPHAEDIT [6]. We demonstrate their performance on CounterFact [19], ZsRE [14], and Wiki-recent [37]. We provide a full description in Appendix B.1.
Evaluation metrics. Following previous work, we evaluate the performance of ME techniques in terms of efficacy, generalization, specificity, and fluency for CounterFact and ZsRE datasets. For Wiki-recent, we additionally evaluate the portability [37] of edited models, which represents the ability to address downstream tasks with edited knowledge. We evaluate the editing efficacy and generalization using the practical metric in Equation 11. When evaluating efficacy on the CounterFact and Wiki-recent datasets, and generalization on CounterFact, we prepend each prompt with 10 distinct 5-token prefixes. Full details of metrics are provided in Section B.2.
Figure 5: PCA visualization of r-k embeddings for 100 edited facts across different models.
6.2 Distinctiveness of Retrieved Embeddings (RQ1)
Finding 1. NAMET promotes more isolated r-k embeddings. We visualize the r-k pairs of 100 facts edited by MEMIT, PMET, and NAMET at layer 4 across models. As shown in Figure 5, the r-k pairs generated by NAMET are more dispersed in the embedding space, while those produced by MEMIT and PMET are more likely to form denser clusters. This observation supports our hypothesis that NAMET encourages more isolated and less entangled target memory embeddings.
@@ 6.3 Performance of Massive Editing (RQ2)
Eff.$\uparrow$ Gen.$\uparrow$ Spe.$\uparrow$ Flu.$\uparrow$ Eff.$\uparrow$ Por.$\uparrow$ Loc.$\uparrow$ Flu.$\uparrow$ Eff.$\uparrow$ Gen.$\uparrow$ Spe.$\uparrow$
LLaMA2-7B FT 0.29 0.23 77.43 4.90 7.23 41.61 36.52 4.91 5.30 4.31 14.69
MEND 0.23 0.31 78.55 3.07 0.00 34.67 37.46 2.69 0.00 0.00 0.50
ROME 0.00 0.00 50.73 4.67 76.73 49.31 51.51 4.97 37.29 6.86 10.27
MEMIT 24.95 22.68 63.84 5.06 34.75 44.93 46.72 5.04 76.63 64.06 15.57
PMET 74.22 46.45 72.47 5.07 81.84 51.11 53.16 4.97 77.29 71.40 16.54
ALPHAEDIT 0.51 0.53 51.14 5.01 0.07 35.34 37.48 5.27 44.26 35.83 12.65
NAMET 89.09 61.21 69.19 5.15 82.51 53.13 54.61 5.08 77.88 73.42 16.65
Qwen2.5-7B FT 16.18 14.15 56.07 5.27 21.17 51.40 51.50 5.15 14.30 13.00 39.28
MEND 0.01 0.06 70.73 2.82 0.00 42.55 44.37 2.72 0.00 0.00 0.09
ROME 0.00 0.00 49.83 5.23 16.28 46.52 46.61 5.02 4.10 3.43 1.30
MEMIT 90.06 63.86 70.53 5.29 94.88 56.97 61.23 5.10 54.12 42.96 31.57
PMET 65.71 52.84 63.14 5.18 82.39 58.38 57.59 5.11 53.58 46.59 36.50
ALPHAEDIT 83.15 55.70 67.16 5.14 94.16 57.17 59.45 5.10 44.52 34.98 25.52
NAMET 90.49 64.37 72.18 5.36 95.61 57.46 60.28 5.09 66.29 60.22 41.80
LLaMA2-13B FT 1.23 0.07 68.57 4.84 13.90 36.89 40.09 4.97 5.95 5.10 16.16
ROME 4.05 1.52 50.44 5.25 11.06 38.14 39.09 4.47 5.52 5.06 2.25
MEMIT 47.98 34.75 71.61 5.17 94.76 51.38 50.40 5.07 69.15 51.58 15.53
PMET 78.60 38.76 81.15 5.26 88.66 49.69 47.58 5.01 53.27 35.73 15.76
ALPHAEDIT 3.03 1.9 54.97 4.21 93.68 51.65 52.33 5.08 80.27 63.66 15.32
NAMET 91.88 61.08 72.01 5.20 95.88 54.08 53.43 5.04 81.24 74.58 15.23
Falcon-7B FT 14.70 13.54 56.34 1.67 23.94 50.46 49.69 3.51 13.64 12.68 32.28
ROME 12.85 12.56 51.48 3.53 74.57 52.10 53.64 5.10 8.39 7.3 10.29
MEMIT 89.21 60.85 77.56 5.19 96.04 55.23 56.91 4.97 82.93 68.93 33.64
PMET 77.61 57.03 70.48 5.17 58.03 54.40 54.49 5.00 69.73 60.69 35.34
ALPHAEDIT 87.62 58.32 72.43 5.00 96.22 55.47 58.02 4.93 53.78 40.83 22.60
NAMET 92.18 63.91 73.86 5.21 96.94 57.08 58.58 5.07 83.18 77.55 36.71
Deepseek-7B FT 2.61 2.49 81.43 5.19 18.85 48.90 52.78 5.00 15.00 12.28 39.14
ROME 0.26 0.30 49.82 5.14 0.55 43.17 46.02 4.06 0.81 0.78 0.75
MEMIT 62.11 42.01 78.04 5.12 33.65 52.28 49.05 4.99 57.10 42.58 39.12
PMET 74.52 43.49 79.01 5.14 86.75 57.85 59.93 5.00 76.97 69.22 39.87
ALPHAEDIT 22.51 14.00 59.92 4.79 18.53 48.33 48.74 4.83 73.41 57.09 34.87
NAMET 89.74 59.98 77.73 5.13 95.14 57.43 60.45 4.97 77.62 71.02 39.16
LLaMA3-8B FT 2.68 1.30 58.16 4.34 16.05 47.51 48.84 4.90 11.75 10.48 42.53
ROME 51.01 33.32 64.37 4.91 82.19 54.94 57.74 5.18 7.40 6.82 27.79
MEMIT 93.76 61.98 77.69 5.26 92.63 55.60 58.75 5.27 78.40 71.76 39.21
PMET 77.71 49.41 71.43 5.10 75.81 56.89 58.26 5.13 68.52 62.72 39.35
ALPHAEDIT 58.97 33.02 85.16 5.37 65.36 51.44 53.55 5.16 64.01 57.01 40.82
NAMET 93.87 65.74 75.07 5.33 93.82 57.88 59.48 5.28 79.34 74.73 44.39
@@
We demonstrate the effectiveness of NAMET in massive editing tasks by comparing it with baseline methods across six popular LLMs. Specifically, we simultaneously edit 10,000 factual associations sampled from the CounterFact and ZsRE datasets. For the Wiki-recent dataset, we modify all 1,266 knowledge items. Additional experiments on other LLMs and datasets are provided in Appendix D. As shown in Table 3, our key findings are as follows:
@@
Eff.$\uparrow$ Gen.$\uparrow$ Spe.$\uparrow$ Flu.$\uparrow$
Qwen2.5-7B MEMIT 77.46 54.34 66.23 5.14
NAMET 83.66 55.31 69.49 5.28
Falcon-7B MEMIT 84.60 56.13 75.51 5.13
NAMET 89.55 61.00 68.44 5.16
LLaMA3-8B MEMIT 87.58 54.76 72.65 5.14
NAMET 91.22 62.24 73.43 5.31
Finding 2. NAMET consistently achieves superior editing performance across all datasets and model architectures. Across all evaluated datasets, NAMET demonstrates the highest levels of editing efficacy and generalization. On the CounterFact dataset, it consistently outperforms other methods, particularly on base models such as LLaMA2-7B, LLaMA2-13B, and Deepseek-7B. For example, NAMET achieves 89.09% efficacy and 61.21% generalization on LLaMA2-7B, outperforming the second-best method (PMET) by 15% on both metrics. The gap widens further compared to MEMIT, with improvements of 65% in efficacy and 39% in generalization. Even on more advanced models such as Qwen2.5-7B, Falcon-7B, and LLaMA3-8B, NAMET consistently surpasses all baselines. Furthermore, its advantage becomes more pronounced at larger editing scales. As shown in Table 4, when editing 15,000 knowledge items on Qwen2.5-7B, NAMET achieves 83.66% efficacy, demonstrating a 10% improvement over MEMIT.
Finding 3. NAMET preserves the general abilities of the edited models. In addition to achieving state-of-the-art editing performance, NAMET does not impair the base model’s fluency or reasoning abilities. Across all datasets, NAMET consistently attains among the highest specificity and fluency scores. Notably, on the Wiki-recent dataset, NAMET achieves the best portability performance on most base models, indicating that the edited models retain their ability to reason about downstream knowledge related to the edited facts.

### v4

@@ 4.2 Impact of Embedding Misalignment on Editing Robustness
Model LLaMA2-7B Deepseek-7B
Prefix Lens 0 5 10 50 0 5 10 50
Editing Acc. 98.50% 84.15% 80.35% 77.40% 99.50% 90.35% 84.25% 85.10%
low $\mathcal{A}(i)$ Acc. - 94.00% 91.00% 90.00% - 94.00% 92.00% 91.00%
top $\mathcal{A}(i)$ Acc. - 46.00% 46.00% 45.00% - 55.00% 54.00% 47.00%
@@
Empirical Verification of Corollary 1. We edit 200 CounterFact facts on LLaMA2-7B and Deepseek-7B using MEMIT, and evaluate average editing accuracy with prefix lengths ranging from 5 to 50 tokens. To assess the impact of embedding misalignment, we also compare the 10 items with the highest and lowest misalignment scores under long-prefix conditions.
@@
Table˜2 shows that editing performance degrades when edited facts are evaluated with prefixes. For LLaMA2-7B, accuracy falls from 98.5% to 84.15% with a 5-token prefix, and further to 77.40% with a 50-token prefix. A similar trend is observed for Deepseek-7B, where editing accuracy drops by around 15% under 50-token prefixes. The robustness to prefix perturbations varies markedly between items prone to embedding misalignment and those that are not. Consistent with Corollary 1, items with lower misalignment scores maintain above-average accuracy, whereas highly misaligned items suffer a sharp decline, with accuracy dropping to an average of 48%.
Corollary 2. Massive editing suffers from degraded performance when multiple samples with the same subject are edited simultaneously. In this case, the reconstruction weight on the target $\beta_{ii}$ decreases while cross-weights $\beta_{ij}$ increase, eventually leading to reconstruction failure of $r_{i}$.
Figure 3: Impact of editing same-subject samples. Shaded region indicates shared items.
Empirical Verification of Corollary 2. Figure 3 shows the reconstruction coefficients for two example subjects under different numbers of edits. Although $\beta_{ii}$ remains the dominant coefficient, its value decreases steadily as the number of co-edited samples increases, while off-diagonal coefficients $\beta_{ij}$ grow accordingly. As a result, $\Delta k_{i}$ is no longer primarily aligned with $r_{i}$ but is instead reconstructed as a mixture of other $r_{j}$, making the recovery of the correct target representation increasingly difficult.
This behavior is consistent with the misalignment measure $\mathcal{A}(i)$, as only $\beta_{ij}$ from the same subject as $i$ take relatively large values, while cross-subject weights remain negligible. Therefore, reconstruction is dominated by the same-subject neighborhood. When $\mathcal{A}(i)$ is small, same-subject embeddings are well aligned across both $k$-space and $r$-space. Thus, using other $r_{j}$ from the same subject to approximate $r_{i}$ introduces only limited error. However, when $\mathcal{A}(i)$ is large, misalignment within this neighborhood amplifies the effect of weight redistribution, causing the residual $\|e_{i}\|$ to grow and ultimately leading to degraded editing performance. We provide details in Appendix˜C.
These findings underscore the strong connection between embedding misalignment and the effectiveness as well as robustness of massive editing. Motivated by this observation, the following section introduces our approach for aligning key and residual embeddings to enhance the overall performance of massive editing.
5 Embedding Alignment Memory Optimization
Figure 4: Method Overview of EAMET.
Motivated by these results, we propose EAMET, which optimizes memory embeddings to promote alignment with key embeddings across facts. This design enhances the model’s ability to edit multiple facts concurrently under practical metrics, while also improving robustness against prefix perturbations and simultaneous edits of same-subject samples. We elaborate on the details below.
Key Embedding Preparation (Figure˜4 (a)). Before optimization, we extract the key embeddings corresponding to each knowledge item that is scheduled for editing. For a given knowledge item $i$, we calculate the cosine similarity between its key embedding $k_{i}$ and the key embeddings of all other items. We then collect these similarity values into the set $P_{k}^{(i)}=\{P_{k}^{(i,j)}=\cos(k_{i},k_{j})\mid j\neq i\}$.
Aligning Memory Embeddings with Key Embeddings (Figure˜4 (b)). For $N$ knowledge items, we separately optimize the target memory embeddings to update factual associations. During the iterative optimization process, we save every optimized residual embedding. When optimizing the target memory for the $i$-th knowledge item, we compute the cosine similarity between $r_{i}$ and all residual embeddings saved so far, and collect them as $P_{r}^{(i)}=\{P_{r}^{(i,j)}\mid j<i\}$. To promote alignment between key and residual embeddings, we compute the KL divergence Chen et al. (2020); He et al. (2020); Sun and Saenko (2016) between $P_{r}^{(i)}$ and $\bar{P}_{k}^{(i)}$, where $\bar{P}_{k}^{(i)}=\{P_{k}^{(i,j)}\mid j<i\}$ denotes the subset of $P_{k}^{(i)}$ corresponding to earlier items:
$L_{\text{KL}}(i)=\text{KL}\!\left(P_{r}^{(i)}\,\|\,\bar{P}_{k}^{(i)}\right).$ (12)
Since KL divergence emphasizes distributional differences, we further strengthen the alignment by selecting the top $M$ cosine similarities $\{P_{k}^{(i,j)}\}$ from $P_{k}^{(i)}$, and computing the mean squared error (MSE) loss between the corresponding residual similarities $\{P_{r}^{(i,j)}\}$:
$L_{\text{MSE}}(i)=\frac{1}{M}\sum_{j=1}^{M}\big\|P_{r}^{(i,j)}-P_{k}^{(i,j)}\big\|^{2}.$ (13)
Optimizing the Target Memory (Figure˜4 (c)). Our goal in this step is to compute the residual update vector $r_{i}$ for each factual association $(s_{i},rel_{i},o_{i})$ such that the model reliably predicts the target object $o_{i}$ while preserving the alignment between the memory embeddings and the key embeddings. To make the optimization procedure explicit, we describe each component of the objective in Equation˜14. For each fact $i$, let $h_{i}^{L}$ denote the hidden state at layer $L$ produced by the templated prompt $tp(s_{i},rel_{i})$. Following prior work (Meng et al., 2022; 2023), we augment this prompt with a set of $N_{\mathrm{FP}}$ randomly sampled prefixes $\{f_{j}\}_{j=1}^{N_{\mathrm{FP}}}$, forming inputs $f_{j}\oplus tp(s_{i},rel_{i})$. These prefixes encourage the model to learn more generalizable memory representations. We write the forward pass of the model with the edited hidden state as $G_{(h_{i}^{L}+=r_{i})}$, indicating that the hidden representation at layer $L$ is perturbed by the update vector $r_{i}$.
Given these definitions, we optimize $r_{i}$ by minimizing the following loss:
$r_{i}=\underset{r_{i}}{\arg\min}\Bigg(\frac{1}{N_{\mathrm{FP}}}\sum_{j=1}^{N_{\mathrm{FP}}}-\log\mathds{P}_{G(h_{i}^{L}+=r_{i})}\!\left[o_{i}\,\middle|\,f_{j}\oplus tp(s_{i},rel_{i})\right]+\lambda_{\mathrm{KL}}L_{\mathrm{KL}}(i)+\lambda_{\mathrm{MSE}}L_{\mathrm{MSE}}(i)\Bigg).$ (14)
Here, the first term encourages the model to predict the correct target object $o_{i}$ under all sampled prefixes. The losses $L_{\mathrm{KL}}(i)$ and $L_{\mathrm{MSE}}(i)$ ensure that the alignment between the memory embeddings and the key embeddings is preserved, with $\lambda_{\mathrm{KL}}$ and $\lambda_{\mathrm{MSE}}$ controlling their relative importance.
The full optimization procedure is detailed in Appendix˜E. We justify our design of combining KL loss and MSE loss in Section˜F.5. As the optimization process is iterative, the editing order of knowledge items may influence the performance of EAMET. We further investigate the robustness of EAMET against different editing orders in Table˜6.
@@ 6 Experiments
In this section, we empirically focus on evaluating the following research questions (RQs). We first demonstrate the effectiveness of EAMET in massive by considering:
  • RQ1. Can EAMET generate more aligned embeddings for different knowledge items?
  • RQ2. How does EAMET perform on massive editing tasks compared with baselines for various LLMs? Can it excel under the practical metric?
We then examine the robustness of EAMET in two representative scenarios:
  • RQ3. How does EAMET perform when evaluating edited facts with prefixes?
  • RQ4. How does EAMET perform when editing multiple facts of the same subject?
@@ 6.1 Experiments Setup
Models, Datasets, and Baselines. We conduct extensive experiments on various LLMs, including LLaMA2-7B (Touvron et al., 2023a), LLaMA2-13B (Touvron et al., 2023a), Falcon-7B (Almazrouei et al., 2023), Qwen-2.5-7B (Yang et al., 2024a), Deepseek-base-7B (Bi et al., 2024), and LLaMA3-8B (Touvron et al., 2023b). We provide additional evaluations on more LLMs in Section˜F.3. We consider a range of ME techniques as baselines: FT (Zhu et al., 2020), MEND (Mitchell et al., 2022a), ROME (Meng et al., 2022), MEMIT (Meng et al., 2023), PMET (Li et al., 2025), and ALPHAEDIT (Fang et al., 2024). We demonstrate their performance on CounterFact (Meng et al., 2022), ZsRE (Levy et al., 2017), and Wiki-recent (Zhang et al., 2024). We provide a full description in Section˜D.1.
Evaluation Metrics. Following previous work, we evaluate the performance of ME techniques in terms of efficacy (Eff.), generalization (Gen.), specificity (Spe.), and fluency (Flu.) for CounterFact and ZsRE datasets. For Wiki-recent, we additionally evaluate the portability (Zhang et al., 2024) (Por.) of edited models, which represents the ability to address downstream tasks with edited knowledge. We propose to evaluate the editing performance of ME techniques by requiring the edited models to strictly examine whether explicit target objects are retrieved, as demonstrated in Figure˜1. The editing efficacy is then defined as:
$\text{Eff.}=\mathds{E}_{i}[o_{i}=\arg\mathop{\max}_{o}\mathds{P}_{f_{\theta}}(o\ |\ (s_{i},rel_{i}))].$ (15)
When evaluating efficacy on the CounterFact and Wiki-recent datasets, and generalization on CounterFact, we prepend each prompt with 10 distinct 5-token prefixes. Full details of metrics are provided in Section˜D.2. We also provide the implementation details of EAMET in Section˜D.3.
Table 3: Misalignment score comparison between different methods. Here, “CF” and “ZS” denote the CounterFact and ZsRE datasets, respectively.
Model EAMET MEMIT PMET
CF ZS CF ZS CF ZS
LLaMA2-7B 377 165 11506 22245 11475 11477
Qwen-7B 374 180 18498 23699 18471 18463
Deepseek-7B 520 161 12135 23241 12155 12046
Falcon-7B 385 181 8564 17589 8602 8590
6.2 Alignment of Retrieved Embeddings (RQ1)
Finding 1. EAMET Promotes More Aligned Embeddings. We compute the summation of the misalignment score between the residual and key embeddings for 10,000 facts edited by MEMIT, PMET, and EAMET under various LLMs. As shown in Table˜3, the residual embeddings generated by EAMET are more aligned with the key embeddings, while those produced by MEMIT and PMET are more likely to cause inconsistency in the key and residual embeddings space. This observation supports our hypothesis that EAMET encourages more aligned target memory embeddings.
@@ 6.3 Performance of Massive Editing (RQ2)
[…]
LLaMA2-7B FT 0.29 0.23 77.43 490.34 7.23 41.61 36.52 491.83 5.30 4.31 14.69
MEND 0.23 0.31 78.55 307.26 0.00 34.67 37.46 269.52 0.00 0.00 0.50
ROME 0.00 0.00 50.73 467.76 76.73 49.31 51.51 497.53 37.29 6.86 10.27
MEMIT 24.95 22.68 63.84 506.69 34.75 44.93 46.72 504.18 76.63 64.06 15.57
PMET 74.22 46.45 72.47 507.10 81.84 51.11 53.16 497.49 77.29 71.40 16.54
ALPHAEDIT 0.51 0.53 51.14 501.63 0.07 35.34 37.48 527.83 44.26 35.83 12.65
EAMET 89.09 61.21 72.19 519.89 93.23 53.13 54.61 503.52 89.47 81.34 15.70
Qwen2.5-7B FT 16.18 14.15 56.07 527.56 21.17 51.40 51.50 515.87 14.30 13.00 39.28
MEND 0.01 0.06 70.73 282.92 0.00 42.55 44.37 272.90 0.00 0.00 0.09
ROME 0.00 0.00 49.83 523.45 16.28 46.52 46.61 502.37 4.10 3.43 1.30
MEMIT 90.06 63.86 70.53 529.27 94.88 56.97 61.23 510.43 54.12 42.96 31.57
PMET 65.71 52.84 63.14 518.92 82.39 58.38 57.59 511.62 53.58 46.59 36.50
ALPHAEDIT 83.15 55.70 67.16 514.07 94.16 57.17 59.45 510.32 44.52 34.98 25.52
EAMET 90.49 64.37 72.18 536.67 95.61 57.46 60.28 509.06 91.03 84.80 41.20
LLaMA2-13B FT 1.23 0.07 68.57 484.56 13.90 36.89 40.09 497.21 5.95 5.10 15.16
ROME 4.05 1.52 50.44 525.12 11.06 38.14 39.09 447.42 5.52 5.06 2.25
MEMIT 47.98 34.75 71.61 517.63 94.76 51.38 50.40 507.84 69.15 51.58 15.53
PMET 78.60 38.76 81.15 526.82 88.66 49.69 47.58 501.61 53.27 35.73 15.76
ALPHAEDIT 3.03 1.9 54.97 421.97 93.68 51.65 52.33 508.82 80.27 63.66 15.32
EAMET 92.85 60.08 77.51 530.78 95.88 52.08 53.43 504.06 87.09 74.58 15.90
Falcon-7B FT 14.70 13.54 56.34 167.18 23.94 50.46 49.69 351.18 13.64 12.68 32.28
ROME 12.85 12.56 51.48 353.38 74.57 52.10 53.64 510.92 8.39 7.3 10.29
MEMIT 89.21 60.85 77.56 519.92 96.04 55.23 56.91 497.35 82.93 68.93 33.64
PMET 77.61 57.03 70.48 517.09 58.03 54.40 54.49 500.87 69.73 60.69 35.34
ALPHAEDIT 87.62 58.32 72.43 500.35 96.22 55.47 58.02 493.56 53.78 40.83 22.60
EAMET 92.37 63.91 78.94 528.98 96.94 57.08 58.58 507.56 92.38 81.15 36.71
Deepseek-7B FT 2.61 2.49 81.43 519.35 18.85 48.90 52.78 500.86 15.00 12.28 39.14
ROME 0.26 0.30 49.82 514.72 0.55 43.17 46.02 406.64 0.81 0.78 0.75
MEMIT 62.11 42.01 78.04 512.16 33.65 52.28 49.05 499.49 57.10 42.58 39.12
PMET 74.52 43.49 79.01 514.58 86.75 57.85 59.93 500.50 76.97 69.22 38.47
ALPHAEDIT 22.51 14.00 59.92 479.52 18.53 48.33 48.74 483.38 73.41 57.09 34.87
EAMET 89.74 59.98 77.73 513.93 97.15 56.43 60.45 501.09 87.27 70.02 39.87
LLaMA3-8B FT 2.68 1.30 58.16 434.67 16.05 47.51 48.84 490.71 11.75 10.48 40.53
ROME 51.01 33.32 64.37 491.98 82.19 54.94 57.74 518.89 7.40 6.82 27.79
MEMIT 93.76 61.98 77.69 526.47 92.63 55.60 58.75 527.29 78.40 71.76 39.21
PMET 77.71 49.41 71.43 510.82 75.81 56.89 58.26 513.84 68.52 62.72 39.35
ALPHAEDIT 58.97 33.02 85.16 537.91 65.36 51.44 53.55 516.53 64.01 57.01 40.82
EAMET 93.87 63.74 79.07 533.30 94.36 57.88 59.48 528.23 85.68 81.34 42.39
@@
We demonstrate the effectiveness of EAMET in massive editing tasks by comparing it with baseline methods across six popular LLMs. Specifically, we simultaneously edit 10,000 factual associations sampled from the CounterFact and ZsRE datasets. For the Wiki-recent dataset, we modify all 1,266 knowledge items. As shown in Table˜4, our key findings are as follows:
@@
[…]
Qwen2.5-7B MEMIT 77.46 54.34 66.23 514.81
EAMET 83.66 55.31 69.49 528.28
Falcon-7B MEMIT 84.60 56.13 75.51 513.82
EAMET 89.55 61.00 68.44 516.91
LLaMA3-8B MEMIT 87.58 54.76 72.65 514.07
EAMET 91.22 62.24 73.43 531.76
Finding 2. EAMET Consistently Achieves Superior Editing Performance Across All Datasets and Model Architectures. Across all evaluated datasets, EAMET demonstrates the highest levels of editing efficacy and generalization. On the CounterFact dataset, it consistently outperforms other methods, particularly on base models such as LLaMA2-7B, LLaMA2-13B, and Deepseek-7B. For example, EAMET achieves 89.09% efficacy and 61.21% generalization on LLaMA2-7B, outperforming the second-best method (PMET) by 15% on both metrics. The gap widens further compared to MEMIT, with improvements of 65% in efficacy and 39% in generalization. Even on more advanced models such as Qwen2.5-7B, Falcon-7B, and LLaMA3-8B, EAMET consistently surpasses all baselines. Furthermore, its advantage becomes more pronounced at larger editing scales. As shown in Table˜5, when editing 15,000 knowledge items on Qwen2.5-7B, EAMET achieves 83.66% efficacy, demonstrating a 10% improvement over MEMIT. We additionally report the superior performance of EAMET across diverse semantic scenarios in Section˜F.1.
Finding 3. EAMET Preserves the General Abilities of the Edited models. In addition to achieving state-of-the-art editing performance, EAMET does not impair the base model’s fluency or reasoning abilities. Across all datasets, EAMET consistently attains among the highest specificity and fluency scores. Notably, on the Wiki-recent dataset, EAMET achieves the best portability performance on most base models, indicating that the edited models retain their ability to reason about downstream knowledge related to the edited facts. We also evaluate the general abilities of edited models on GLUE (Wang et al., 2018) and find that EAMET yields minimal deviation from pre-edit performance (Section˜F.2).
Table 6: Impact of editing sequence on EAMET’s performance on Counterfact and ZsRE datasets.
Method Counterfact ZsRE
[…]
EAMET (original sequence) 89.09 61.21 72.19 519.06 89.47 81.34 15.70
– random shuffle (seed=0) 88.21 60.79 71.84 519.21 87.63 77.42 15.56
– random shuffle (seed=1) 89.11 60.78 72.03 518.84 86.99 76.08 15.58
– random shuffle (seed=2) 88.91 59.38 72.34 518.23 87.56 77.47 15.59
As EAMET preserves previously optimized residual embeddings when updating new knowledge items, the editing sequence could potentially affect its performance. To assess this, we examine EAMET’s robustness under different editing orders on the Counterfact and ZsRE datasets. In Counterfact, all knowledge items have distinct subjects, whereas in ZsRE some items share the same subject and are adjacent in the original order. We therefore randomly shuffle the order of 10,000 items three times and report the average performance, alongside the original sequence as a reference.
Finding 4. EAMET is Robust to Editing Sequence. As shown in Table˜6, EAMET’s performance remains stable across editing orders. On Counterfact, random shuffles produce only negligible variations in efficacy, generalization, and specificity. On ZsRE, editing efficacy shows a slight decline of about 2%, likely due to the neighborhood structure of items sharing the same subject in the original sequence. Overall, these results suggest that EAMET is largely insensitive to editing order, demonstrating strong robustness to sequence variations.